OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)? y = 2x – 8 y = – 2x + 6 y = 2x – 11 y = – 2x + 9

OpenStudy (anonymous):

y = 2x – 11

OpenStudy (hoblos):

slope = (-1+7)/(5-2) = 6/3 = 2 y+1 = 2(x-5) y= 2x -11

OpenStudy (anonymous):

y = 2x – 11